OpenStudy (anonymous):

Simplify (7 – 2i) – (3 + 9i)

OpenStudy (anonymous):

This is for imaginary numbers? I found a great site that explains everything about them: http://betterexplained.com/articles/a-visual-intuitive-guide-to-imaginary-numbers/

OpenStudy (callisto):

(7 – 2i) – (3 + 9i) = 7 - 2i - 3 - 9i = 7 - 3 -2i - 9i = 7 - 3 - (2+9)i = ...?

OpenStudy (anonymous):

4-11i

OpenStudy (anonymous):

= 4-11i

OpenStudy (callisto):

Yes :)

OpenStudy (anonymous):

can you help me on another one?

OpenStudy (callisto):

My pleasure~

OpenStudy (anonymous):

\[9/(7-3\sqrt{3})-(6-4\sqrt{3})\]

OpenStudy (callisto):

Is this your question? \[\frac{9}{(7-3\sqrt3) - (6-4\sqrt3)} \]